In theory, what will be the ratio of the rate of effusion of potassium to the rate of effusion of germanium?
Answer #1

Rate of effusion is inversely proportional to square root of molar mass

rate = k*sqrt(1/M)

M is molar mass

If we take ratio, above expression becomes

rate(K)/rate(Ge) = sqrt (M(Ge)/M(K))

we have:

M(K) = 39.1 g/mol

M(Ge) = 72.63 g/mol

rate(K)/rate(Ge) = sqrt (72.63/39.1)

rate(K)/rate(Ge) = sqrt (1.8575)

rate(K)/rate(Ge) = 1.3629

Answer: 1.36